OpenStudy (anonymous):

HELP! If x_n=x_n-1+x_n-2 Then what is the limit n->infinty of x_n/x_n-1? I know this will be (1+sqrt5)/2 but how do I prove?

OpenStudy (anonymous):

call the limit \(x\) and solve \[x=1+\frac{1}{x}\]

OpenStudy (anonymous):

Satellite73 can you go back to question

OpenStudy (anonymous):

But if I say the limit is X, what do I do to show it is equal to 1 + 1/X?

OpenStudy (anonymous):

\[\frac{x_n}{x_{n-1}}=\frac{x_{n-1}+x_{n-2}}{x_{n-1}}\] so \[\frac{x_n}{x_{n-1}}=1+\frac{x_{n-1}}{x_{n-2}}\]

OpenStudy (anonymous):

if \[\lim_{n\to \infty}\frac{x_n}{x_{n-1}}=x\] then \[\lim_{n\to \infty}\frac{x_{n-2}}{x_{n-1}}=\frac{1}{x}\] since it is the reciprocal of the first one

OpenStudy (anonymous):

i made a typo above sorry it should be \[\frac{x_n}{x_{n-1}}=1+\frac{x_{n-2}}{x_{n-1}}\]

OpenStudy (anonymous):

I understand everything but where limit x_n-2/x_n-1= 1/x?

OpenStudy (anonymous):

Ohhhh, wait. I understand it now. So then is simplifies to X = 1+1/X and then you times both sides by X and turn it into a quadratic.